2/3(11x - 1/6) - 3/4 = 1/4
Answer by user_dude2008(1862)

2/3(11x - 1/6) - 3/4 = 1/4
22/3x - 1/9 - 3/4 = 1/4
264x - 4 - 27 = 9
264x - 31 = 9
264x = 40
x=40/264
x=5/33
Answer: x=5/33